Find the equation of the line that goes through the points (-4,1) and (-1,4)
Answer by Alan3354(69443)

Find the equation of the line that goes through the points (-4,1) and (-1,4)
Use determinants.
------------
|x +y 1|
|-4 +1 1| = 0
|-1 +4 1|
----------
x(1-4) - y(-4+1) + -16+1 = 0
-3x + 3y = 15
x - y = -5